determine whether the graph of the polynomial has y-axis symmetry, origin symmetry or neither
f(x)=7-x^4

Answer by stanbon(75887)

determine whether the graph of the polynomial has y-axis symmetry, origin symmetry or neither
f(x)=7-x^4
f(-x) = 7 - (-x)^4 = 7 - x^4
-f(-x) = -(7 - x^4) = -7 + x^4
--------------------
Since f(x) = f(-x) the function is symmetric to the y-axis.
========================
Cheers,
Stan H.